OpenStudy (anonymous):

Evaluate : (4.1 x 10^7)(2 x 10^7)

OpenStudy (anonymous):

multiply \[4.1\times 2\] get \[8.2\] add the exponents get \[8.2\times 10^{7+7}=8.2\times 10^{14}\]